A photon of energy E is scattered by a particle of rest energy Eo. Find the maximum kinetic energy of the recoiling particle in terms of E and Eo.
1
Expert's answer
2020-05-05T18:41:54-0400

from crompton scatttering we get that :


"K=hf\frac{\Delta\lambda}{\lambda+\Delta\lambda}"


"K=hf\dfrac{2h}{m_oc(\lambda+\frac{2h}{m_oc})}"


"K=\dfrac{2hfhf}{m
_oc(\lambda f+\frac{2hf}{m_oc})}"


"K=\dfrac{2hfhf}{m
_oc^2(1+\frac{2hf}{m_oc^2})}"


now lets substitute the vale of E and Eo

E=hf and Eo=moc2


"K=\dfrac{2E^2}{E_o(1+\frac{E}{E_o})}"